1. Write the acid dissociation reaction equation for Na2H2PO4.

2. Write a reaction equation showing how the buffer solution maintains pH ~ 8.0 upon addition of acid.

3. Write a reaction equation showing how the buffer solution maintains pH ~ 8.0 upon addition of a base. Buffer solution: Acid= Na2H2PO4 base Na2HPO4

4. Would you expect this buffer to be better at resisting a change in pH upon addition of acid, or upon addition of base? Why?
Answer #1

1. H2PO4-1 (aq) = HPO4-2 (aq) + H+ (aq)

2. When you add acid (H+) in the above equilibrium, the equilibrium shifts to the back ward direction ( left side) to maintain the pH.

HPO4-2 (aq) + H+ (aq) = H2PO4-1 (aq)

3. When you add base, H+ is removed by base and the equilibrium shifts forward direction to maintain the pH.

H2PO4-1 (aq) = HPO4-2 (aq) + H+ (aq)

4. The buffer's pH is at 8.

It is basic pH. If you add acid it will resist the pH change more effectively.